OpenStudy (anonymous):

x^Sqroot log x= 10^8

OpenStudy (anonymous):

\[\large x^{\sqrt{\log(x)}}=10^8\]?

OpenStudy (anonymous):

Yes thats correct

OpenStudy (anonymous):

hmmm is that log base ten?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i guess you could start by taking the log base ten of both sides and get \[\log(x^{\sqrt{\log(x)}})=8\]

OpenStudy (anonymous):

or \[\sqrt{\log(x)}\log(x)=8\]

OpenStudy (anonymous):

yeah ive gotten that far, i just dont know what to do next

OpenStudy (anonymous):

yeah me neither

OpenStudy (anonymous):

Darn I was hoping this site would help :(

OpenStudy (anonymous):

maybe write it as \[\log(x)^{\frac{3}{2}}=8\]

OpenStudy (anonymous):

that should make it easier, since you solve \[u^{\frac{3}{2}}=8\] via \[u=8^{\frac{2}{3}}=2^2=4\]

OpenStudy (anonymous):

making \[\log(x)=4\] which, in equivalent exponential form, is \(x=10^4\)

OpenStudy (anonymous):

i wonder if we could have seen that from the start...

OpenStudy (anonymous):

So... What would be your ending answer?

OpenStudy (anonymous):

that would be it \(10^4\) want to check it?

OpenStudy (anonymous):

yeah thats correct. But you kind of lost me when u said u^3/2=8

OpenStudy (anonymous):

is it clear how i went from \[\sqrt{\log(x)}\log(x)\] to \[\left(\log(x)\right)^{\frac{3}{2}}\]

OpenStudy (anonymous):

No sorry

OpenStudy (anonymous):

i added the exponents \[\sqrt{u}=u^{\frac{1}{2}}\] so \[\sqrt{u}\times u=u^{\frac{3}{2}}\]

OpenStudy (anonymous):

Oh I understand

OpenStudy (anonymous):

So now I am at log x ^3/2 =8

OpenStudy (anonymous):

then to solve for \(\log(x)\) take the cubed root of both sides, you get \[(\log(x))^{\frac{1}{2}}=2\] then square both sides and get \[\log(x)=4\]

OpenStudy (anonymous):

U are a genius, Thank you so much

OpenStudy (anonymous):

Do you get paid to do this?

OpenStudy (anonymous):

i get $10 for every answer $20 for each correct one

OpenStudy (anonymous):

seriously? how do u sign up for that job

OpenStudy (anonymous):

and how do they keep track if you get it right or not?